OpenStudy (iwanttogotostanford):

PLEASE HELP:

OpenStudy (elijahrich1113):

can you write it in numbers?

OpenStudy (iwanttogotostanford):

@elijahrich1113 f(x)=13/10-x

myininaya (myininaya):

Think: When are fractions not defined? What number can we not divide by?

OpenStudy (iwanttogotostanford):

@jim_thompson5910 please help, no one has helped me all day!

jimthompson5910 (jim_thompson5910):

The function is this \[\Large f(x) = \frac{13}{10-x}\] right?

OpenStudy (iwanttogotostanford):

yes

jimthompson5910 (jim_thompson5910):

Recall that you CANNOT divide by zero. You cannot have something like \(\Large \frac{1}{0}\) Zero must NEVER be in the denominator

jimthompson5910 (jim_thompson5910):

The question I have for you @iwanttogotostanford is this: what value of x makes the denominator \(\Large 10-x\) equal to zero?

jimthompson5910 (jim_thompson5910):

so basically I'm asking you to solve \[\Large 10-x = 0\] for x

OpenStudy (iwanttogotostanford):

ok so it would end up to be x= x+10

jimthompson5910 (jim_thompson5910):

if you added x to both sides, you'd end up with \(\Large x = 10\)

OpenStudy (iwanttogotostanford):

sorry i meant the other x to be a 0

jimthompson5910 (jim_thompson5910):

so if \(\Large x = 10\), then \(\Large 10-x\) is equal to 0

jimthompson5910 (jim_thompson5910):

x = 10 causes a division by zero error so we must kick it out of the domain. No other x value gives us this trouble, so we can keep any other real number we want

OpenStudy (iwanttogotostanford):

oh ok i see that

jimthompson5910 (jim_thompson5910):

In set notation, you would say that the domain is \[\Large \left\{x|x\in\mathbb{R} \ , \ x \ne 10\right\}\] which is fancy math speak for "x is any real number but NOT 10. So anything but 10"

OpenStudy (iwanttogotostanford):

so it would be all real numbers except for 10, correct?

jimthompson5910 (jim_thompson5910):

In interval notation, the domain would be written as \[\Large \left(-\infty,10\right)\cup\left(10,\infty\right)\] which is the interval from -infinity to +infinity, but we take out 10. So we form a hole at 10 on the number line and keep everything else.

jimthompson5910 (jim_thompson5910):

`so it would be all real numbers except for 10, correct?` yes @iwanttogotostanford

OpenStudy (iwanttogotostanford):

thank yoU! you are always the best help honestly. Clear and concise help ! I appreciate it. Can i please ask you a few more? You really help me learn the concepts !

jimthompson5910 (jim_thompson5910):

if you have any other questions, post them one at a time in a new thread (one question per post)

jimthompson5910 (jim_thompson5910):

Then tag me in any question you get stuck on